How many grams of CuCl2·2H2O are needed to prepare 0.500 L of solution containing 0.250 M Cl-(aq)?


1
Expert's answer
2022-06-13T02:18:03-0400

CM (Cl-) = 0,25 M

V = 0,5 L

Mr (CuCl2 • 2H2O) = 171 g/mol


n(Cl-) = CM * V = 0,25 * 0,5 = 0,125 mol Cl-


CuCl2 • 2H2O => Cu2+ + 2 Cl- + 2H2O


1mol CuCl2 • 2H2O => 2mol Cl-

X => 0,125mol Cl-


X = 0,125 * 1 / 2 = 0,0625 mol


m = n * Mr

m(CuCl2 • 2H2O) = 0,0625 * 171 = 10,6875 g


Answer: 10,6875 g CuCl2 • 2H2O